x²y"-(2x+1)y'+y²=0


1
Expert's answer
2020-08-18T18:39:46-0400
"Solution"

Write the differential equation in standard form: "y''-\frac{2x+1}{x^2}y'+\frac{y^2}{x^2}=0"


We have "P(x)=-\frac{2x+1}{x^2}" and "Q(x)=\frac{1}{x^2}" , so "x= 0" is a singular point. Taking limits we get "lim_{x\u21920}xP(x) = 1" and "lim_{x\u21920}x^2Q(x) = 0" . Therefore "x= 0" is a regular singular point. We look for a solution of the form "y(x) =\u2211^\u221e_{n=0}a_nx^{n+r}" . Plug the series into the equation to get


"y(x) =∑^∞_{n=0}a_nx^{n+r}\\
 y' =∑^∞_{n=0}(n+r)a_nx^{n+r-1}\\
 y'' =∑^∞_{n=0}(n+r)(n+r-1)a_nx^{n+r-2}"


"y'' =x^2\u2211^\u221e_{n=0}(n+r)(n+r-1)a_nx^{n+r-2}-2x\u2211^\u221e_{n=0}(n+r)a_nx^{n+r-1}-\u2211^\u221e_{n=0}(n+r)a_nx^{n+r-1}+(\u2211^\u221e_{n=0}a_nx^{n+r})^2=0"


"\implies y'' =∑^∞_{n=0}(n+r)(n+r-1)a_nx^{n+r}-∑^∞_{n=0}2(n+r)a_nx^{n+r}-∑^∞_{n=0}(n+r)a_nx^{n+r-1}+(∑^∞_{n=0}a_nx^{n+r})^2=0"


Upon re-arrangement we get


"(r(r-1)-2r)a_0x^r+\sum_{n=0}^\infin \{[(n+r+1)(n+r+3)]q_{n+1}-(n+r-1)a_n\}x^{n+r}"


Equating the powers of x yields



"(r^2-3r)a_0=0,\ \ a_{n+1}=\frac{a_n}{n+r+3}, for\ n=0,\ 1,\ ...."

But "a_0 \ne 0" implies that "r= 0,3" . Solving for the "a_n\u2019s" we get

For r=0

"a_{n+1}=\frac{a_n}{n+3}\\
a_1=\frac{a_0}{3}\\
a_2=\frac{a_1}{4}=\frac{a_0}{4 \cdot 3}=\frac{a_0 \cdot 2 \cdot 1}{4 \cdot 3 \cdot 2 \cdot 1}=\frac{2!}{4!}a_0 \\ 
a_3=\frac{a_2}{5}=\frac{a_0}{5 \cdot 4 \cdot 3}=\frac{a_0 \cdot 2 \cdot 1}{5 \cdot4 \cdot 3 \cdot 2 \cdot 1}=\frac{2!}{5!}a_0 \\
.\\
.\\
.\\
a_n=\frac{2!}{(n+3)!}a_0"


For r=3

"a_{n+1}=\frac{a_n}{n+6}\\
a_1=\frac{a_0}{6}\\
a_2=\frac{a_1}{7}=\frac{a_0}{7 \cdot 6}=\frac{a_0}{7 \cdot 6!}\\ 
a_3=\frac{a_2}{8}=\frac{a_0}{8 \cdot 7 \cdot 6}=\frac{a_0}{8 \cdot7 \cdot 6!} \\
.\\
.\\
.\\
a_n=\frac{a_0}{(n+6)!}"


Two linearly independent Frobenius series (with "a_0= 1" ) are given by



"y_1(x)=\sum_{n=0}^\infin a_nx^n=\sum_{n=0}^\infin \frac{2!}{(n+3)!}x_n,\ y_2(x)=\sum_{n=0}^\infin a_nx^{n+3}=x^3\sum_{n=0}^\infin \frac{x_n}{(n+6)!}"